Take this same regression equation (y = 12 + 4.5x) and calculate the residual when x and y have the following actual values x = 8 and y = 52. The residual for this particular point is:

52

48

4

-4
Answer #1

Answer: 4

The regression equation: y = 12 + 4.5x

Actual value of y when x = 8 is 52

Predicted value of y when x = 8 is (12 + 4.5*8) = 48

The residual for this particular point is (Actual – Predicted) = (52 – 48) = 4